OpenStudy (anonymous):

A 84.0-kg person stands on a scale in an elevator. What is the apparent weight when the elevator is (a) accelerating upward with an acceleration of 1.40 m/s2, N (b) moving upward at a constant speed, and N (c) accelerating downward with an acceleration of 1.50 m/s2?

OpenStudy (anonymous):

alright so i dont have a calculator with me so i'm going to just write down the equations. a) F=ma = 84*(9.8+1.40) m = F/9.8 b) only acceleration will affect the weight constant speed in any direction will not. so therefore m=84kg c) same as par (a) only this time since the elevator is moving down, you will feel much lighter as you are going against gravity, so just minus 1.5m/s/s from 9.8m/s/s so F=ma=84*(9.8-1.50) m=F/9.8 done

OpenStudy (anonymous):

makes sense? if not let me know, i can try to explain more.

OpenStudy (anonymous):

Thank you so much and yes it does! :)

OpenStudy (anonymous):

glad i could help.